OpenStudy (ja1):

Ok I went blank, what do you do:

OpenStudy (ja1):

\[0 = -16x^{2} + 73.16x + 5.9\]

OpenStudy (ja1):

There is no GCF so what do I do?

OpenStudy (tpaulus):

Quadratic formula... http://www.wolframalpha.com/input/?i=0%3D%E2%88%9216x2%2B73.16x%2B5.9 x = 4.65177,-0.0792709

OpenStudy (ja1):

Yes but I do not need the answer, besides the fact that it is against the CoC I need to know what to do.

OpenStudy (ja1):

I don't need the answer, I need to know how to do this.

OpenStudy (ja1):

Besides giving away answers is against the CoC: http://openstudy.com/code-of-conduct

OpenStudy (uri):

Lol! \(\huge\sf Hey,~you~want~Ansurs?\)

OpenStudy (espex):

If you are finding roots, plug your coefficients into the quadratic formula. \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (uri):

Yes put the Quadratic formula yo! Use your calculator. :)

OpenStudy (espex):

Or by hand, like a BEAST! :P

OpenStudy (uri):

Lol it's in points,so i CAN never use my brain..I need a calculator.

OpenStudy (uri):

@JA1 I think you know how to do it.

OpenStudy (ja1):

Oh sorry I was afk, anyways I need to factor not use the quadratic formula.

OpenStudy (ja1):

@eSpeX

OpenStudy (uri):

Use middle term breaking method.

OpenStudy (ja1):

Ok I think I know what it means but refresh me?

OpenStudy (uri):

You have to break the middle term in such a way that when you add them it becomes the middle term and when you multiply them it becomes the last term Example, \[x ^{2}+7x-7\] x^2 +8x -1x-7 So when we add or subract both of them it comes 7x and when we multiply them it becomes -7x.

OpenStudy (ja1):

Ok so how would I do that for this?

OpenStudy (uri):

There is no common,but wait let me think :D

OpenStudy (ja1):

xD kk

OpenStudy (ja1):

Ah I'll just cheat ;)

OpenStudy (uri):

Wait,I'm solving!

OpenStudy (ja1):

lol kk

OpenStudy (uri):

Ughh I'm getting a totally different answer.